OpenStudy (anonymous):

Two lines form a linear system. The slope of one line is 4 and it has a y-intercept of (0, -3). The slope of the other line is 4 and it has a y-intercept of (0, 8). How many solutions does this system have?

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (acxbox22):

y=mx+b m=slope and b=y intercept

OpenStudy (yanasidlinskiy):

Just like on the other one I helped you with. What do you think?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

I thought it was 2

OpenStudy (yanasidlinskiy):

http://www.mathops.com/free/a1ls008.php Here's a website!!:) No. Can you solve for it and tell me what you got?

OpenStudy (anonymous):

1

OpenStudy (yanasidlinskiy):

I'm not sure myself. Lemme solve it!!:) Lol.

OpenStudy (anonymous):

ok

OpenStudy (yanasidlinskiy):

For some reason..That's not what I'm getting. Can you post your steps?

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

wait what did u get

OpenStudy (yanasidlinskiy):

-3

OpenStudy (anonymous):

here are the answers 0 1 2 infinitely many

OpenStudy (yanasidlinskiy):

Oh..Hahhaha!! Wait..

OpenStudy (anonymous):

what

OpenStudy (yanasidlinskiy):

What did you plug in for the formula? How did you do it?

OpenStudy (anonymous):

I guessed sorry iam not good a t math can u explain to me how to do it

OpenStudy (yanasidlinskiy):

Hahha!! Ok!:) I'll definitely will:) http://www.math.com/school/subject2/lessons/S2U4L2GL.html Does that look easy to you? ^

OpenStudy (yanasidlinskiy):

Do you understand how to do it now?

OpenStudy (anonymous):

it gave me a head acke

OpenStudy (anonymous):

could u explain it

OpenStudy (yanasidlinskiy):

Yes! y=mx+b y=4(0)-3 y=0+3 y=-3 Does that make sense? Can you show your steps for the other equation? After that, we'll determine what's the answer:)

OpenStudy (anonymous):